# Extension principles for the Einstein Yang–Mills system

### v1

@@ 3.1.4. Main energy estimates
$\iint_{\mathcal{D}}\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}\cdot\ell_{1}\mathrm{d}u\mathrm{d}\underline{u}=-\iint_{\mathcal{D}}\left(1-\mu\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$ (3.17)
[…]
@@
$\iint_{\mathcal{D}}\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}\cdot\ell_{i}\,\mathrm{d}u\mathrm{d}\underline{u}$
@@
$\lesssim$ $\left(\iint_{\mathcal{D}}\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}\cdot\left(\iint_{\mathcal{D}}\Omega^{2}\left(\frac{\alpha^{2}}{r^{2+b}}+\frac{\Phi^{2}}{r^{2+b}}\right)\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}$
@@
is not allowed, since we have no yet any improved bounds for the energies on the right hand side of the inequality above. The difficulty coming from these quadratic terms is resolved via applying integration by parts. In this way, we expect to eventually gain spacetime integrals, such as $\iint_{\mathcal{D}}\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$, $\iint_{\mathcal{D}}\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$, with favourable signs.
@@
Among all the terms on the right side of (3.1.4), the nonlinear terms in the second and third lines can be in principle estimated by means of (after multiplying $r^{-1-b}\cdot\Omega^{2}h\alpha^{\prime}$)
@@
$\lesssim\varepsilon\Lambda\iint_{\mathcal{D}}\left(\frac{\Phi^{2}}{r^{2+b}}+\frac{\alpha^{2}}{r^{2+b}}+\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.18)
@@
The extra smallness of (3.18) suggests they could be absorbed, since in the preceding steps of treating the linear terms $\ell_{i}$, we have managed to extract a positive bulk terms as $\iint_{\mathcal{D}}\left(\frac{\Phi^{2}}{r^{2+b}}+\frac{\alpha^{2}}{r^{2+b}}+\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$ on the left hand side of the energy inequality through integration by parts.
@@
We now go on with the detailed proof. First of all, let us turn to the left hand side of (3.1.4). After multiplied with $\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}$, it becomes,
$\partial_{u}\alpha^{\prime}\cdot\Omega^{2}h\frac{\alpha^{\prime}}{r^{1+b}}$
[…]
We notice that the last term
@@
$\frac{1+b}{2}\underline{h}h\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2+b}}=-\frac{1+b}{2}(1-\mu)\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2+b}}$
@@
which indeed lies on the left hand side the energy identity has a bad sign. Nevertheless, this term is compensated by (3.17) and their summation still contributes a positive term
@@
$\iint_{\mathcal{D}}\frac{1-b}{2}\left(1-\mu\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u},\quad 0\leq b<\frac{\sqrt{89}-9}{4},$
on the left side of the energy identity.
[…]
$r^{-1-b}\Omega^{2}h\alpha^{\prime}\cdot\ell_{2}=-w^{2}h^{-1}r^{-1-b}\partial_{\underline{u}}\alpha^{2}$
[…]
$+w^{2}r^{-1-b}\alpha^{2}\partial_{\underline{u}}h^{-1}+w^{2}\partial_{\underline{u}}r^{-1-b}\alpha^{2}h^{-1}$
[…]
$+2w^{2}\Omega^{2}r^{-2-b}\alpha^{4}-(1+b)w^{2}\Omega^{2}r^{-2-b}\alpha^{2},$
@@
where we have made use of the equation (2.4a) in the last identity. In particular, we notice that the last term $-(1+b)w^{2}\Omega^{2}r^{-2-b}\alpha^{2}$ which lies on the right hand side the energy identity admits a favourable sign. When $i=3$, we have
$r^{-1-b}\Omega^{2}h\alpha^{\prime}\cdot\ell_{3}=2h^{-1}r^{-2-b}(w^{2}-1)w\cdot\partial_{\underline{u}}\alpha$
[…]
$-\partial_{\underline{u}}w\cdot 2r^{-2-b}h^{-1}(w^{2}-1)\alpha$
[…]
$-4h^{-2}\Omega^{-2}\frac{(\partial_{\underline{u}}w)^{2}}{r^{3+b}}(w^{2}-1)w\alpha-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}-6(w^{2}-1)\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}.$
@@
Here the quadratic term $-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}$ admits a favourable sign, while the sign of $(4+2b)\Omega^{2}\frac{w^{2}-1}{r^{3+b}}w\alpha$ is not definite. Nevertheless, we can perform the same trick on this quadratic term as follows
$(4+2b)\Omega^{2}\frac{w^{2}-1}{r^{3+b}}w\alpha=(2+b)h^{-1}\frac{w^{2}-1}{r^{3+b}}\partial_{\underline{u}}(w^{2}-1)$
[…]
$+(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{(w^{2}-1)^{2}}{r^{4+b}}.$
@@
$r^{-1-b}\Omega^{2}h\alpha^{\prime}\cdot\ell_{3}={}$ $\partial_{\underline{u}}\left(2wh^{-1}\frac{\Phi\alpha}{r^{1+b}}\right)+\left(1+\frac{b}{2}\right)\partial_{\underline{u}}\left(h^{-1}\frac{\Phi^{2}}{r^{1+b}}\right)$
$-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}+(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}$
$-2\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}\Phi^{2}-4\Omega^{2}\frac{\Phi}{r^{2+b}}w\alpha^{3}-6\Omega^{2}\Phi\frac{\alpha^{2}}{r^{1+b}},$
where in the second line, the quadratic term $-4\Omega^{2}\frac{\alpha^{2}}{r^{2+b}}$ contributes a positive term in the energy identity, while $(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}$ offers a bad sign. Nevertheless, noting the positive integral $\iint_{\mathcal{D}}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left hand side of the first order energy identity (3.1.4), the bad-sign term $(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}$ will be absorbed when (3.1.4) is added into the second order energy identity.
Taking all the calculations into accounts, we integrate these identities over $\iint_{\mathcal{D}}\mathrm{d}u\mathrm{d}\underline{u}$ to derive the energy identity
$\int_{C_{u}}\frac{(\alpha^{\prime})^{2}}{2r^{1+b}}\Omega^{2}h\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\frac{1}{r^{1+b}}\left(\left(w\alpha-\Phi\right)^{2}-\left(2+\frac{b}{2}\right)\Phi^{2}\right)h^{-1}\mathrm{d}u$
$+\iint_{\mathcal{D}}\left(\frac{1-b}{2}+\frac{3\Phi^{2}}{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left((5+b)+(2+b)\Phi^{2}\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$-\iint_{\mathcal{D}}(3+b)\left(1+\frac{b}{2}\right)\frac{\Phi^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{D}}\left(2(1-\Phi^{2})\alpha^{2}+4\Phi^{2}-2\mu-6w\Phi\alpha-r\Phi\right)\frac{\alpha\alpha^{\prime}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(2\left(wr\alpha+w^{2}\alpha^{2}\right)-4w\Phi\alpha-(7+b)r\Phi\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.19)
Now in view of the identities (3.1.4), (3.1.4) and (3.1.4), to eliminate the negative term $-\iint_{\mathcal{D}}(3+b)\left(1+\frac{b}{2}\right)\Omega^{2}\frac{\Phi^{2}}{r^{2+b}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left hand side of (3.1.4), and obtain coercive energy on $\underline{C}_{\underline{u}}$, we consider the linear combination
@@
$\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left((9+2b)\alpha^{2}+(\alpha^{\prime})^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
$+\iint_{\mathcal{D}}\left((1-b)+3\Phi^{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left((1-9b-2b^{2})+2(2+b)\Phi^{2}+(9b+2b^{2})\mu\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(\frac{3(3+b)}{2}+(4+b)\mu\right)\frac{\Phi^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(2(1+b)+4\alpha^{2}\right)\frac{h^{2}(\partial_{u}w)^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
[…]
$+\iint_{\mathcal{D}}\left(\mu+r|\Phi|+\Phi^{2}+\alpha^{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(\mu+r|\Phi|+\Phi^{2}+\alpha^{2}+r|\alpha|\right)\frac{\alpha^{2}}{r^{2+b}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.20)
We note that, since $0\leq b<\frac{\sqrt{89}-9}{4}$ and hence $01-9b-2b^{2}>$ 0, terms on the left hand side of (3.1.4) are all positive. In particular, the energy on $\underline{C}_{\underline{u}}$ is coercive (the energy on $C_{u}$ is obviously coercive), and there is
$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
$\lesssim$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u.$
In view of the bootstrap assumption (3.10) and the fact that $\mu$ is small, and $r$ is finite in $\mathcal{D}$, the spacetime integrals on the right hand side of (3.1.4) can be absorbed by the bulk integrals on the left hand side. We thus achieve the energy bound
$\int_{C_{u}}\frac{\Omega^{2}h}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+(\alpha^{\prime})^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r^{1+b}}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u\lesssim I_{1},$ (3.21)
and then (3.8) follows.
3.1.5. Pointwise estimates
With the help of the energy estimates obtained in the previous step, we will deduce various $L^{\infty}$ estimates to improve the bootstrap assumption.

### v2

@@ 3.5. Main energy estimates
$\iint_{\mathcal{D}}\Omega^{2}h\frac{\alpha^{\prime}}{r}\cdot\ell_{1}\mathrm{d}u\mathrm{d}\underline{u}=-\iint_{\mathcal{D}}\left(1-\mu\right)\frac{(\alpha^{\prime})^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$ (3.15)
admits a favourable sign and can be moved to the left-hand side. In contrast, the remaining terms $\iint_{\mathcal{D}}r^{-1}\cdot\Omega^{2}h\alpha^{\prime}\cdot\ell_{i}\,\mathrm{d}u\mathrm{d}\underline{u}$, i=2, 3i=2,\,3, which are now quadratic, lack definite signs. A straightforward estimate such as
@@
$\iint_{\mathcal{D}}\Omega^{2}h\frac{\alpha^{\prime}}{r}\cdot\ell_{i}\,\mathrm{d}u\mathrm{d}\underline{u}$
@@
$\lesssim$ $\left(\iint_{\mathcal{D}}\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}\cdot\left(\iint_{\mathcal{D}}\Omega^{2}\left(\frac{\alpha^{2}}{r^{2}}+\frac{\Phi^{2}}{r^{2}}\right)\mathrm{d}u\mathrm{d}\underline{u}\right)^{\frac{1}{2}}$
@@
is not allowed at this stage, as we do not yet have improved bounds for the spacetime integrals on the right-hand side. The difficulty posed by these quadratic terms is resolved via applying integration by parts. This allows us to ultimately obtain spacetime integrals with favorable signs, such as $\iint_{\mathcal{D}}\Omega^{2}\frac{\alpha^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$, $\iint_{\mathcal{D}}\Omega^{2}\frac{\Phi^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$.
@@
Among all the terms on the right-hand side, the nonlinear terms in the second and third lines can, in principle, be estimated by means of (after multiplication by $r^{-1}\cdot\Omega^{2}h\alpha^{\prime}$)
@@
$\lesssim\varepsilon\Lambda\iint_{\mathcal{D}}\left(\frac{\Phi^{2}}{r^{2}}+\frac{\alpha^{2}}{r^{2}}+\frac{(\alpha^{\prime})^{2}}{r^{2}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.16)
@@
Actually, the term (3.16) is expected to be absorbed, since during the treatment of the linear terms $\ell_{i}$, we have already managed to generate positive bulk terms of the form $\iint_{\mathcal{D}}\left(\frac{\Phi^{2}}{r^{2}}+\frac{\alpha^{2}}{r^{2}}+\frac{(\alpha^{\prime})^{2}}{r^{2}}\right)\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$ on the left-hand side of the energy inequality via integration by parts.
@@
We now go on with the detailed proof. First of all, we analyze the left-hand side of (3.5). After multiplied $\Omega^{2}h\frac{\alpha^{\prime}}{r}$, it becomes,
$\partial_{u}\alpha^{\prime}\cdot\Omega^{2}h\frac{\alpha^{\prime}}{r}$
$={}$ $\partial_{u}\left(\Omega^{2}h\frac{(\alpha^{\prime})^{2}}{2r}\right)-\partial_{u}\left(\Omega^{2}h\right)\frac{(\alpha^{\prime})^{2}}{2r}-\frac{1}{2}\partial_{u}\left(\frac{1}{r}\right)\Omega^{2}h(\alpha^{\prime})^{2}$
@@
$={}$ $\partial_{u}\left(\Omega^{2}h\frac{(\alpha^{\prime})^{2}}{2r}\right)+\Omega^{2}\left(\frac{\mu}{r}-\frac{Q^{2}}{r^{3}}\right)\frac{(\alpha^{\prime})^{2}}{2r}+\frac{1}{2}\underline{h}h\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2}}.$
We observe that the last term
$\frac{1}{2}\underline{h}h\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2}}=-\frac{1}{2}(1-\mu)\Omega^{2}\frac{(\alpha^{\prime})^{2}}{r^{2}}$
@@
on the left-hand side of the energy identity carries an unfavorable sign. Nevertheless, this term can be absorbed by the positive contribution from (3.15), and ultimately yields a net positive term
@@
$\iint_{\mathcal{D}}\frac{1}{2}\left(1-\mu\right)\frac{(\alpha^{\prime})^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
on the left-hand side of the energy identity.
To proceed, we focus on the two terms $\iint_{\mathcal{D}}r^{-1}\Omega^{2}h\alpha^{\prime}\cdot\ell_{i}\,\mathrm{d}u\mathrm{d}\underline{u}$, i=2, 3i=2,\,3. When $i=2$, it holds that
$r^{-1}\Omega^{2}h\alpha^{\prime}\cdot\ell_{2}=-w^{2}h^{-1}r^{-1}\partial_{\underline{u}}\alpha^{2}$
$={}$ $-\partial_{\underline{u}}\left(w^{2}h^{-1}r^{-1}\alpha^{2}\right)+\partial_{\underline{u}}w^{2}h^{-1}r^{-1}\alpha^{2}$
$+w^{2}r^{-1}\alpha^{2}\partial_{\underline{u}}h^{-1}+w^{2}\partial_{\underline{u}}r^{-1}\alpha^{2}h^{-1}$
$={}$ $-\partial_{\underline{u}}\left(w^{2}h^{-1}r^{-1}\alpha^{2}\right)+2\Omega^{2}wr^{-1}\alpha^{3}$
$+2w^{2}\Omega^{2}r^{-2}\alpha^{4}-w^{2}\Omega^{2}r^{-2}\alpha^{2},$
@@
where we have made use of the equation (2.4a) in the last identity. In particular, we note that the last term $w^{2}\Omega^{2}r^{-2}\alpha^{2}$ on the right-hand side the energy identity has a favourable sign. When $i=3$, we have
$r^{-1}\Omega^{2}h\alpha^{\prime}\cdot\ell_{3}=2h^{-1}r^{-2}(w^{2}-1)w\cdot\partial_{\underline{u}}\alpha$
$={}$ $\partial_{\underline{u}}\left(2r^{-2}h^{-1}(w^{2}-1)w\alpha\right)-\partial_{\underline{u}}\left(2r^{-2}h^{-1}(w^{2}-1)\right)w\alpha$
$-\partial_{\underline{u}}w\cdot\left(2r^{-2}h^{-1}(w^{2}-1)\alpha\right)$
$={}$ $\partial_{\underline{u}}\left(2r^{-2}h^{-1}(w^{2}-1)w\alpha\right)+4\Omega^{2}r^{-3}(w^{2}-1)w\alpha$
$-4h^{-2}\Omega^{-2}\frac{(\partial_{\underline{u}}w)^{2}}{r^{3}}(w^{2}-1)w\alpha-4\Omega^{2}\frac{\alpha^{2}}{r^{2}}-6(w^{2}-1)\Omega^{2}\frac{\alpha^{2}}{r^{2}}.$
@@
Here the quadratic term $4\Omega^{2}\frac{\alpha^{2}}{r^{2}}$ admits a favourable sign, whereas $4\Omega^{2}\frac{w^{2}-1}{r^{3}}w\alpha$ lacks a definite sign. Nevertheless, we can perform the same technique to this quadratic term to proceed
$4\Omega^{2}\frac{w^{2}-1}{r^{3}}w\alpha=2h^{-1}\frac{w^{2}-1}{r^{3}}\partial_{\underline{u}}(w^{2}-1)$
$={}$ $\partial_{\underline{u}}\left(h^{-1}\frac{(w^{2}-1)^{2}}{r^{3}}\right)-\partial_{\underline{u}}h^{-1}\frac{(w^{2}-1)^{2}}{r^{3}}+3\Omega^{2}\frac{(w^{2}-1)^{2}}{r^{4}}.$
@@
$r^{-1}\Omega^{2}h\alpha^{\prime}\cdot\ell_{3}$
$={}$ $\partial_{\underline{u}}\left(2wh^{-1}\frac{\Phi\alpha}{r}\right)+\partial_{\underline{u}}\left(h^{-1}\frac{\Phi^{2}}{r}\right)$
$-4\Omega^{2}\frac{\alpha^{2}}{r^{2}}+3\Omega^{2}\frac{\Phi^{2}}{r^{2}}-2\Omega^{2}\frac{\alpha^{2}}{r^{2}}\Phi^{2}-4\Omega^{2}\frac{\Phi}{r^{2}}w\alpha^{3}-6\Omega^{2}\Phi\frac{\alpha^{2}}{r}.$
In the third line, the quadratic term $4\Omega^{2}\frac{\alpha^{2}}{r^{2}}$ contributes positively to the energy identity, while $3\Omega^{2}\frac{\Phi^{2}}{r^{2}}$ has an unfavorable sign. Nevertheless, noting the presence of the positive integral $\iint_{\mathcal{D}}\frac{\Phi^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left-hand side of the first order energy identity (3.5), the term $3\Omega^{2}\frac{\Phi^{2}}{r^{2}}$ can be absorbed when (3.5) is taken into accounts.
Integrating over $\iint_{\mathcal{D}}\mathrm{d}u\mathrm{d}\underline{u}$ leads to the energy identity
$\int_{C_{u}}\frac{(\alpha^{\prime})^{2}}{2r}\Omega^{2}h\mathrm{d}\underline{u}+\int_{\underline{C}_{\underline{u}}}\frac{1}{r}\left(\left(w\alpha-\Phi\right)^{2}-2\Phi^{2}\right)h^{-1}\mathrm{d}u$
$+\iint_{\mathcal{D}}\left(\frac{1}{2}+\frac{3\Phi^{2}}{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(5+2\Phi^{2}\right)\frac{\alpha^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}-\iint_{\mathcal{D}}3\frac{\Phi^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$={}$ $\int_{C_{u_{0}}}\frac{(\alpha^{\prime})^{2}}{2r}\Omega^{2}h\mathrm{d}\underline{u}+\iint_{\mathcal{D}}2\mu\frac{(\alpha^{\prime})^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(2(1-\Phi^{2})\alpha^{2}+4\Phi^{2}-2\mu-6w\Phi\alpha-r\Phi\right)\frac{\alpha\alpha^{\prime}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(2\left(wr\alpha+w^{2}\alpha^{2}\right)-4w\Phi\alpha-7r\Phi\right)\frac{\alpha^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.17)
Now in view of the identities (3.5), (3.5) and (3.5), we aim to eliminate the negative term $\iint_{\mathcal{D}}3\Omega^{2}\frac{\Phi^{2}}{r^{2}}\mathrm{d}u\mathrm{d}\underline{u}$ on the left-hand side of (3.5), and meanwhile obtain a coercive energy on $\underline{C}_{\underline{u}}$. To achieve this, we form the linear combination
@@
$\int_{C_{u}}\frac{\Omega^{2}h}{r}\left(9\alpha^{2}+(\alpha^{\prime})^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
$+\iint_{\mathcal{D}}\left(1+3\Phi^{2}\right)\frac{(\alpha^{\prime})^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}+\iint_{\mathcal{D}}\left(1+4\Phi^{2}\right)\frac{\alpha^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}\left(\frac{9}{2}+4\mu\right)\frac{\Phi^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}+\iint_{\mathcal{D}}\left(2+4\alpha^{2}\right)\frac{h^{2}(\partial_{u}w)^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$={}$ $\int_{C_{u_{0}}}\frac{\Omega^{2}h}{r}\left(9\alpha^{2}+(\alpha^{\prime})^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}+\iint_{\mathcal{D}}\Phi^{2}\frac{\Phi^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}2\left(2(1-\Phi^{2})\alpha^{2}+4\Phi^{2}-2\mu-6w\Phi\alpha-r\Phi\right)\frac{\alpha\alpha^{\prime}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}$
$+\iint_{\mathcal{D}}2\left(2\left(wr\alpha+w^{2}\alpha^{2}\right)-4w\Phi\alpha-7r\Phi\right)\frac{\alpha^{2}}{r^{2}}\Omega^{2}\mathrm{d}u\mathrm{d}\underline{u}.$ (3.18)
We note that, terms on the left-hand side of (3.5) are all positive. In particular, the energy on $\underline{C}_{\underline{u}}$ is coercive (the energy on $C_{u}$ is obviously coercive), and there is
$\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u$
$\lesssim$ $\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(2\left(w\alpha-\Phi\right)^{2}+\frac{\Phi^{2}}{2}+2h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u.$
In view of the bootstrap assumption (3.8), the smallness of $\mu$, and the boundedness of $r$ in $\mathcal{D}$, the spacetime integrals on the right-hand side of (3.5) can be absorbed by the bulk integrals on the left-hand side. We thus achieve the energy bound
$\int_{C_{u}}\frac{\Omega^{2}h}{r}\left(\alpha^{2}+(\alpha^{\prime})^{2}+\Phi^{2}\right)\mathrm{d}\underline{u}$
$+\int_{\underline{C}_{\underline{u}}}\frac{h^{-1}}{r}\left(\Phi^{2}+\alpha^{2}+h^{2}(\partial_{u}w)^{2}\right)\mathrm{d}u\lesssim I_{1}.$ (3.19)
As a result, the energy bound (3.6) follows.
3.6. Pointwise estimates
With the help of energy estimates derived earlier, we now proceed to establish various $L^{\infty}$ estimates in order to improve the bootstrap assumption.